OpenStudy (anonymous):

Part 1: Fill in the missing row of the truth table. Part 2: Are the two statements logically equivalent? Why or why not? (~p ∧ q) and (~p ∨ q)

OpenStudy (anonymous):

p q ~ p ~p ∧ q ~p∨ q (~p ∧ q) ↔ (~p ∨ q) T T F F T F T F F F F T F F T F T F

OpenStudy (rsadhvika):

hint - first two columns of third row are : F T

OpenStudy (anonymous):

IS it. FTTFFT?

OpenStudy (rsadhvika):

no

OpenStudy (rsadhvika):

~p ^ q = T ^ T = T

OpenStudy (rsadhvika):

so it starts with FTTT

OpenStudy (anonymous):

So is it FTTTFT?